Hi, What is the derivation of formula for sum of sine/cosine series, the angles of which are in A.P.? Thanks, Nishtha

Hi,


What is the derivation of formula for sum of sine/cosine series, the angles of which are in A.P.?


Thanks,


Nishtha

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

 

Dear nishtha,

sin ( π + θ ) = –sinθ and sin (2π + θ ) = sinθ

                                       ...  –sin (α + β) = sin(π +  )

... trigonometric-series

  so on. Using these results, required sum is

  S=sin α+ trigonometric-series… to n terms

    = sin α + sin trigonometric-series... to n terms trigonometric-series

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

sumesh
9 years ago
if x1, x2, x3......xn in AP then prove that
secx1.secx2 + secx2.secx3 + secx3.secx4................................secxn-1.secxn = (tanxn-tanx1)/sin(x2-x1)